# The physical basis of teleological evolutions

### v12

@@ 2 The teleological notion in the quantum domain
This section provides the physical basis to the notion of teleological evolution in the quantum domain. It is in two parts. In the first, we formalize the notion of completing the ordinary quantum description of quantum algorithm by time-symmetrizing it. This is what highlights the teleological character of quantum algorithms. The second is the generalization from quantum algorithms to quantum correlation.
@@ 2.1 Completing the ordinary quantum description of quantum algorithms by time-symmetrizing it
(1) Extend the usual quantum description of the quantum algorithm, limited to the process of solving the problem, to the upstream process of setting the problem. We call this extended quantum description the ordinary quantum description of the quantum algorithm: ordinary because it is still a causal description that cannot describe causal loops. We will work on the simplest, four-drawer instance of Grover’s quantum search algorithm and then generalize it.
(2) This point is propaedeutic to point (4) below. We show how to relativize with respect to the problem-solver the ordinary quantum description of the quantum algorithm, which is with respect to the customary observer external to the quantum process. Of course, the problem-solver must be concealed from the problem-setting selected by the initial measurement: it would give her full information about the solution of the problem555 This is obvious in Grover algorithm, where the setting and the solution of the problem are both the number of the drawer with the ball. More in general, we must assume that the problem-solver does not know the problem-setting selected by the initial measurement but knows everything about the structure of the problem, thus also the relation between problem-setting and solution. .
@@
(4) Relativize with respect to the problem-solver the causal loops generated by the time-symmetrizing the ordinary quantum description of the quantum algorithm, which are of course with respect to the customary external observer. Note that the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup, each telling the problem-solver in advance one of the possible halves of the information about the solution, are of course with respect to her.
@@ 2.1.1 Extending the usual quantum description of quantum algorithms to the process of setting the problem
Let us introduce the notation. We make reference to the simplest four drawer instance of Grover algorithm:
@@
$|00\rangle_{A}$ and $00$; $|01\rangle_{A}$ and $01$; $|10\rangle_{A}$ and $10$; $|11\rangle_{A}$ and $11$;
we call it content of $A$ observable. Measuring $\hat{A}$ at the end of the quantum algorithm yields the number of the drawer with the ball selected by the initial measurement, namely the solution of the problem. Note that $\hat{B}$ and $\hat{A}$ commute.
@@
We should start from the quantum state in the upper-left corner of Table I diagram and then follow the vertical and horizontal arrows. Initially, at time $t_{1}$, both registers, each by itself, are in a quantum superposition of all their basis vectors. By measuring $\hat{B}$ (the content of $B$) in this initial state, Bob selects the number of the drawer with the ball at random, say it comes out $01$ – see the state of register $B$ under the vertical arrow666 By the way, Bob would be free to unitarily change at will the sorted out number of the drawer with the ball. We omit this operation that would change nothing here. .
@@
Note that we will never need to know $\hat{\mathbf{U}}_{1,2}$ (i.e. Grover algorithm). We only need to know that there exist a unitary transformation between the outcome of the initial measurement and that of the final measurement, what is always the case since the final measurement outcome is an invertible function of the initial one.
@@
Note that this relativization relies on relational quantum mechanics $\left[15,16\right]$ where the quantum state is relative to the observer (also note that one observer is external and the other internal to the quantum process). In the following, we will continue to call the quantum description with respect to the external observer the ordinary quantum description of the quantum algorithm. Instead we will always specify that the other quantum description is that with respect to the problem-solver.
@@ 2.1.3 Time-symmetrizing the ordinary quantum description of the quantum algorithm
Note that the selection, by the initial measurement, of half of the information that specifies the problem-setting is made possible by two facts: (i) the measurement occurs in a quantum superposition of all the possible problem settings, therefore it is possible to reduce it to the measurement of any partition of all the possible problem settings and (ii) the problem-setting never changes along the unitary evolution between the two measurement outcomes, therefore the measurement of the partition of problem-settings can be postponed along that evolution until making it simultaneous with the final measurement. Of course, as the even sharing of the selection can be performed in many possible mutually exclusive ways, eventually we should take their quantum superposition.
@@
The initial measurement of $\hat{B}_{l}$, selecting the 00 of $01$, projects the initial quantum superposition of all the basis vectors of register $B$ on the superposition of those beginning with 00 – vertical arrow on the left of the diagram. The latter superposition causally propagates forward in time, by $\hat{\mathbf{U}}_{1,2}$, into the superposition of the two tensor products on the right of the right looking horizontal arrows. Then the final measurement of $\hat{A_{r}}$, selecting the $1$ of $01$, projects the latter superposition on the term ending with $1$ under the right vertical arrow. Note that this outcome of the final measurement, which inherits both selections, coincides with that of the ordinary quantum description – see the bottom-right corner of Table I diagram. Time-symmetrically, this final measurement outcome causally propagates backward in time, by $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$, until becoming the definitive outcome of the initial measurement. This backward in time propagation (which inherits both selections) is an instance of the time-symmetrization of the ordinary quantum description of the quantum algorithm – bottom line of the diagram. Note that, mathematically (disregarding the time-direction of causality) it is the ordinary quantum description of the unitary part of the quantum algorithm back again see the bottom line of the diagram in Table I. Indeed, the symbol “$\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow$” is mathematically equivalent to “$\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow$”.
@@ 2.1.4 Relativizing the causal loops with respect to the problem-solver
The causal loops generated by time-symmetrizing the ordinary quantum description of the quantum algorithm, which is with respect to the external observer, do not yet map on those logically implied by the sheer existence of the quantum speedup, which are with respect to the problem-solver. We should then relativize the cuasal loops of Table III with respect to the problem-solver. We should just conceal from her the part of the problem-setting selected by the initial measurement. This is done by postponing after her problem-solving action the projection of the quantum state due to the initial measurement of $\hat{B}_{l}$ This yields the causal loop in the following table:
@@
Also note that, for the mathematical equivalence between $\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow$ and $\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow$, the bottom line of the diagram of Table IV can also be read in the usual left to right way:
@@ 2.2 Generalization
First, one can see that everything immediately generalizes to any number of drawers: it suffices to increase in an even way the number of quantum bits in each quantum register.
@@
The second observation is that we never needed to look inside $\hat{\mathbf{U}}_{1,2}$. This unitary transformation is just defined by the fact that it must build up one-to-one correlation between two initially unrelated commuting observables without changing the eigenvalue of the observable selected by the initial measurement, see Table I. Besides Grover algorithm, we could be dealing with any quantum algorithm where the solution is an invertible function of the problem-setting and is determined with certainty. By the way, if the solution is not an invertible function of the problem-setting, it suffices that, at the end, the problem-solver, before reading the solution, measures the problem-setting. In any such quantum algorithm, the problem-solver always knows in advance half of the information about the solution she will produce and measure in the future and can use this knowledge to produce the solution with fewer computation steps. In the optimal quantum case, the number of computation steps is that needed to solve the problem in an optimal logical way benefiting of the advanced knowledge in question. Of course, this explains the quadratic quantum speedup of Grover quantum search algorithm, which solves an unstructured problem. In $\left[7\right]$, we have seen that, when the problem is structured, the advanced knowledge of half of the information about the solution can create the room for an exponential speedup.
@@
Things generalize also beyond quantum algorithms. Let us consider a unitary evolution between an initial and a final one-to-one correlated observables that does not change the eigenvalue of the observable selected by the initial measurement. In this kind of process, one is free to postpone the initial measurement at will. When made simultaneous with the final measurement, the time-symmetrization we are dealing with must already be in place to become the perfect atemporal symmetry between the simultaneous measurements of two maximally entangled observables – see the states above and below the vertical arrow in Table II. This is for the principle of sufficient reason: there would be no reason for the prevalence of either measurement in the simultaneous selection of the pair of one-to-one correlated measurement outcomes.
We can therefore, more in general, state that:
@@
The ordinary quantum description of any unitary evolution between an initial and final one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement is incomplete and must be completed by time-symmetrizing it.
@@
(i) The unitary evolution builds up one-to-one correlation between two initially unrelated observables. This is, of course, the case of quantum algorithms but it can also be that of processes of identical form (as that Table I but for a higher number of quantum bits) that are not quantum algorithms. In any case the building up of the quantum correlation can occur with quantum speedup. If the process is not a quantum algorithm, the notion of advanced knowledge, on the part of the problem-solver, of half of the information about the solution she will produce and measure in the future might loose any meaning. However, it can be replaced by the notion of an equivalent reduction of the dimension of the Hilbert space in which the search for one-to-one correlation takes place.
(ii) The unitary evolution does not build up any quantum correlation because the two observables are maximally entangled to start with. We have seen in the Rationale that this is the case of quantum nonlocality.
@@
The notion that computation is possible in the quantum world and the fundamental notion of quantum bit was introduced by Finkelstein $[19]$ in 1969; that was in the search for a unification between quantum mechanics and relativity theory, more specifically in the context of seeing the universe as a quantum computer. In 1982, Feynman $\left[20\right]$ highlighted the higher than classical computational efficiency of the quantum world, noticing that the classical simulation of a quantum process can require an amount of physical resources exponentially higher than the amount involved in the quantum process itself. In 1985, Deutsch $\left[21\right]$ devised the seminal quantum algorithm that delivers a quantum computational speedup.
@@ 3 The possible physical basis of the teleological evolutions of life
We have seen that quantum algorithms are full fledged teleological evolutions. This raises the question whether the physical basis of their teleological character could also apply to the natural evolutions for which the teleological notion was conceived.
Of course, a difficulty is seeing quantum features in evolutions of life that can have the size of the living system of Earth and have durations of billions of years. We have seen in the Rationale that the only possible way out is deriving these features from an assumed quantum evolution of the universe.
A consequent quantum interpretation of the Fine-tuned Universe version of the Anthropic Principle would say that the evolution of the universe toward the “value” of life (in particular, “existence of sentient life” or “absence of life”) under the fundamental physical laws and constants would build up a quantum correlation between the value of the fundamental constants since the beginning of time and the value of life today. The evolution in question could therefore occur with quantum speedup, being consequently teleological in character. In the universe we are in, of course, the same would apply to the evolution of the universe toward life and then – even if decohered – to the evolution of life on Earth and consequently to the evolutions of species. Like in the case of random quantum search algorithms $\left[14\right]$, under the (conservative) assumption that the search for one-to-one correlation (for the value of life) is random, the room for quantum speedup would become an actual quantum speedup.
In the following subsections, after encoding a toy model of the evolution of the universe toward the value of life on the quantum description of Grover quantum search algorithm, we examine the impact that a teleological character of the evolution of the living would have on different areas of science and knowledge.
@@ 3.1 Teleological evolution of the universe toward the value of life – a toy model
We should imagine a value of the fundamental constants observable $\hat{B}$ and a value of life observable $\hat{A}$. Say that the eigenvalue $01$ of $\hat{B}$ is the value of the fundamental constants compatible with the development of sentient life and the eigenvalues $00,10,11$ the values of those incompatible. We assume that, correspondingly, selecting the eigenvalue $01$ by the measurement of $\hat{A}$ at the present time would mean existence of sentient life; selecting the eigenvalues $00,10,11$ would mean instead absence of life.
@@
The universe’s unitary evolution under the fundamental physical laws and constants would have finally, at the present time, built up a quantum superposition of tensor products, each an eigenstate of $\hat{B}$ multiplying the corresponding eigenstate of $\hat{A}$ – see Table II. In our universe, at the present time one with observers, the measurement of $\hat{A}$ at the present time would select the eigenvalue $01$ of $\hat{A}$, namely “existence of sentient life” and the corresponding eigenvalue $01$ of the value of the fundamental physical constants observable $\hat{B}$. Since the universe’s evolution does not change the value of the fundamental constants, it is as if the eigenvalue of $\hat{B}$ was selected by a measurement of $\hat{B}$ at the beginning of time, when there could be no observers.
Of course, this is exactly the quantum search algorithm described in the previous section, the version with respect to the problem-solver of Table II. As shown by its time-symmetrization (Table IV), it can undergo with a quantum speedup and does so if the quantum search for the solution is random $\left[14\right]$.

### v13

@@ 3 The teleological notion in the quantum domain
This section provides the physical basis of the notion of teleological evolution in the quantum domain. It is in two parts. In the first, we formalize the notion of completing the ordinary quantum description of quantum algorithm by time-symmetrizing it. The second is the generalization from quantum algorithms to quantum correlation.
@@ 3.1 Completing the ordinary quantum description of quantum algorithms by time-symmetrizing it
(1) Extend the usual quantum description of the quantum algorithm, limited to the process of solving the problem, to the upstream process of setting the problem. We will work on the simplest, four-drawer instance of Grover’s quantum search algorithm and then generalize.
(2) The ordinary quantum description of the quantum algorithm is with respect to the customary observer external to the quantum process. We show how to relativize it with respect to the problem solver, to whom the problem setting selected by the initial measure is to be kept hidden.
@@
(4) Relativize with respect to the problem-solver the causal loops generated by time-symmetrizing the ordinary quantum description of the quantum algorithm, which are of course with respect to the customary external observer. Note that the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup, each telling the problem-solver in advance one of the possible halves of the information about the solution, are of course with respect to her.
Steps (1) to (4) are developed in the following subsections.
@@ 3.1.1 Extending the usual quantum description of quantum algorithms to the process of setting the problem
Let us introduce the notation. We make reference to the simplest, four drawer, instance of Grover algorithm:
@@
$|00\rangle_{A}$ and $00$; $|01\rangle_{A}$ and $01$; $|10\rangle_{A}$ and $10$; $|11\rangle_{A}$ and $11$.
Measuring $\hat{A}$ at the end of the quantum algorithm yields the number of the drawer with the ball selected by the initial measurement, namely the solution of the problem. Note that $\hat{B}$ and $\hat{A}$ commute.
@@
We should start from the quantum state in the upper-left corner of Table I diagram and then follow the vertical and horizontal arrows. Initially, at time $t_{1}$, both registers, each by itself, are in a quantum superposition of all their basis vectors. By measuring $\hat{B}$ (the content of $B$) in this initial state, Bob selects the number of the drawer with the ball at random, say it comes out $01$ – see the state of register $B$ under the vertical arrow555 By the way, Bob would be free to unitarily change at will the sorted out number of the drawer with the ball. We omit this operation that would change nothing here. .
@@
Note that we will never need to know $\hat{\mathbf{U}}_{1,2}$ (i.e. Grover algorithm). We only need to know that there can be a unitary transformation between the outcome of the initial measurement and that of the final measurement, what is always the case since the final measurement outcome is an invertible function of the initial one.
@@
Note that this relativization relies on relational quantum mechanics $\left[16,17\right]$ where the quantum state is relative to the observer (also note that, here, one observer is external and the other internal to the quantum process). In the following, we will continue to call the quantum description with respect to the external observer the ordinary quantum description of the quantum algorithm. Instead we will always specify that the other quantum description is that with respect to the problem-solver.
@@ 3.1.3 Time-symmetrizing the ordinary quantum description of the quantum algorithm
Note that the selection, by the initial measurement, of half of the information that specifies the problem-setting is made possible by two facts: (i) the measurement occurs in a quantum superposition of all the possible problem-setting, therefore it is sufficient to reduce it to the measurement of one of the possible halves of the bits of the problem-setting in question (in register $B$) and (ii) the problem-setting never changes along the unitary evolution between the two measurement outcomes, therefore the measurement of the other half of the bits can be postponed along that evolution until making it simultaneous with the final measurement. Of course, as the even sharing of the selection can be performed in many possible mutually exclusive ways, eventually we should take their quantum superposition.
@@
The initial measurement of $\hat{B}_{l}$, selecting the 00 of $01$, projects the initial quantum superposition of all the basis vectors of register $B$ on the superposition of those beginning with 00 – vertical arrow on the left of the diagram. The latter superposition causally propagates forward in time, by $\hat{\mathbf{U}}_{1,2}$, into the superposition of the two tensor products on the right of the right looking horizontal arrows. Then the final measurement of $\hat{A_{r}}$, selecting the $1$ of $01$, projects the latter superposition on the term ending with $1$ under the right vertical arrow. Note that this outcome of the final measurement, which inherits both selections, coincides with that of the ordinary quantum description of the quantum algorithm – see the bottom-right corner of Table I diagram. Time-symmetrically, this final measurement outcome causally propagates backward in time, by $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$, until becoming the definitive outcome of the initial measurement. This backward in time propagation (which inherits both selections) is an instance of the time-symmetrization of the ordinary quantum description of the quantum algorithm – bottom line of the diagram. Note that, mathematically (disregarding the time-direction of causality) it is the ordinary quantum description of the unitary part of the quantum algorithm back again – see the bottom line of the diagram in Table I. Indeed, the symbol “$\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow$” is mathematically equivalent to “$\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow$”.
@@ 3.1.4 Relativizing the causal loops with respect to the problem-solver
The causal loops generated by time-symmetrizing the ordinary quantum description of the quantum algorithm, which is with respect to the external observer, do not yet map on those logically implied by the sheer existence of the quantum speedup, which are with respect to the problem-solver. Indeed, we should relativize the causal loops of Table III with respect to the problem-solver. We should just conceal from her the part of the problem-setting selected by the initial measurement. This is done by postponing after her problem-solving action the projection of the quantum state due to the initial measurement of $\hat{B}_{l}$ This yields the causal loop in the following table:
@@
Also note that, for the mathematical equivalence between $\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow$ and $\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow$, the bottom line of Table IV diagram (mathematically) can also be read in the usual left to right way:
@@ 3.2 Generalization
First, everything immediately generalizes to any number, $N$, of drawers: it suffices to increase from $2=\sqrt{4}$ to $\sqrt{N}$ the number of quantum bits in each quantum register. We can see that all considerations remain unaltered; Alice (the problem-solver) always knows in advance one of the possible halves of the information about the solution and can use this knowledge to produce the solution with fewer computation steps, $\operatorname*{O}\left(\sqrt{N}\right)$ in the optimal quantum case. In equivalent terms, the dimension of the Hilbert space in which the search for the solution (or for one-to-one correlation between the two observables) takes place reduces from $N$ basis vectors to $\sqrt{N}$ basis vectors.
@@
Second, we never needed to look inside $\hat{\mathbf{U}}_{1,2}$. This unitary transformation is just defined by the fact that it must build up one-to-one correlation between two initially unrelated commuting observables ($\hat{B}$ and $\hat{A}$) without changing the eigenvalue of the observable ($\hat{B}$) selected by the initial measurement, like in Table I in the case of $2$ quantum bits per register. Besides Grover algorithm, we could be dealing with any quantum algorithm where the solution is an invertible function of the problem-setting and is determined with certainty. In any such quantum algorithm, the problem-solver always knows in advance half of the information about the solution she will produce and measure in the future and can use this knowledge to produce the solution with fewer computation steps. In the optimal quantum case, the number of computation steps is that needed to solve the problem in an optimal logical way benefiting of the advanced knowledge in question. Of course, this explains the quadratic quantum speedup of Grover quantum search algorithm, where the search is in an unstructured domain. In $\left[4\right]$, we have seen that, when the search is in a structured domain, the advanced knowledge of half of the information about the solution can create the room for an exponential speedup.
@@
Things generalize also beyond quantum algorithms. Let us consider a unitary evolution between an initial and a final one-to-one correlated observables that does not change the eigenvalue of the observable selected by the initial measurement. In this kind of process, one is free to postpone the initial measurement at will. When made simultaneous with the final measurement, the time-symmetrization we are dealing with must already be in place to become the perfect atemporal symmetry between the simultaneous measurements of two maximally entangled observables – see the states above and below the vertical arrow in Table II, thinking that the measurement of $\hat{B}$ is postponed and made simultaneous with that of $\hat{A}$. This is for the principle of sufficient reason: there would be no reason for the prevalence of either measurement in the simultaneous selection of the pair of one-to-one correlated measurement outcomes.
Note that, in the case of quantum algorithms, the above reason for time-symmetrizing is redundant with the previous reason that that is necessary being the only way of generating the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup. Since the time-symmetrization we are dealing with changes our way of viewing physical reality, the redundancy in question should not be despised.
For the more general reason, we can state that:
@@
The ordinary quantum description of any unitary evolution between an initial and a final one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement is incomplete and is completed by time-symmetrizing it.
@@
(i) The unitary evolution builds up one-to-one correlation between two initially unrelated observables. This is, of course, the case of quantum algorithms but it can also be that any process of identical form. We mean the form of Table I in the case of the two quantum bits per register. Note that all the considerations of the previous subsection apply and the building up of the quantum correlation can always occur with an at least quadratic quantum speedup. Of course, if the process is not a quantum algorithm, the notion of advanced knowledge, on the part of the problem-solver, of half of the information about the solution she will produce and measure in the future must be replaced by the notion of an equivalent reduction of the dimension of the Hilbert space (from $N$ basis vectors to $\sqrt{N}$ basis vectors) in which the search for one-to-one correlation takes place.
(ii) The unitary evolution does not build up any quantum correlation because the two observables are maximally entangled to start with. We have seen in the Rationale that this is the case of quantum nonlocality. Here, the causal loops generated by the time-symmetrization play the role of the hidden variables envisaged by Einstein and the others to locally explain quantum nonlocality.
@@
The notion that computation is possible in the quantum world and the fundamental notion of quantum bit was introduced by Finkelstein $[20]$ in 1969; that was in the context of seeing the universe as a quantum computer (a cellular automaton), while searching for a unification between quantum mechanics and relativity theory. In 1982, Feynman $\left[21\right]$ highlighted the higher than classical computational efficiency of the quantum world, noticing that the classical simulation of a quantum process can require an amount of physical resources exponentially higher than the amount involved in the quantum process itself. In 1985, Deutsch $\left[22\right]$ devised the seminal quantum algorithm that delivers a quantum computational speedup.
@@ 4 The possible physical basis of the teleological evolutions of life
The fact that quantum algorithms are full fledged teleological evolutions raises the question whether the physical basis of their teleological character could also apply to the natural evolutions to which the teleological notion was once applied – is sometimes applied also today.
We have seen in the Rationale that the only possible way of seeing quantum features in the evolutions of life, which can be the size of Earth and have durations of billions of years, is deriving them from an assumed quantum evolution of the universe.
A consequent quantum interpretation of the Fine-tuned Universe version of the Anthropic Principle says that the evolution of the universe toward the “value” of life (in particular, “existence of sentient life” or “absence of life”) under the fundamental physical laws and constants builds up a one-to-one quantum correlation between the value of the fundamental constants since the beginning of time and the value of life today. The evolution in question can therefore occur with quantum speedup and consequently be teleological in character. The “can occur” is in the sense that there is certainly the room for its occurring with quantum speedup. Like in the case of random quantum search algorithms $\left[15\right]$, under the (conservative) assumption that the search for this one-to-one correlation is random, the room for quantum speedup becomes an actual quantum speedup. In the universe we are in, of course, the same would apply to the evolution of the universe toward life and, cascading, to the decohered evolution of life on Earth and the evolutions of species.
In the following subsections, after simply encoding a toy model of the evolution of the universe toward the value of life on the quantum description of Grover quantum search algorithm, we examine the impact that a teleological character of the evolution of the living would have on different areas of science and knowledge.
@@ 4.1 Teleological evolution of the universe toward the value of life – a toy model
We should imagine a value of the fundamental constants observable $\hat{B}$ and a value of life observable $\hat{A}$. Say that the eigenvalue $01$ of $\hat{B}$ is the value of the fundamental constants compatible with the development of sentient life and the eigenvalues $00,10,11$ the values of those incompatible. Correspondingly, let the eigenvalue $01$ of $\hat{A}$ mean existence of sentient life; the eigenvalues $00,10,11$ absence of life.
@@
The universe’s unitary evolution under the fundamental physical laws and constants would have finally, at the present time, built up a quantum superposition of tensor products, each an eigenstate of $\hat{B}$ multiplying the corresponding eigenstate of $\hat{A}$ – see Table II. In the universe we are in, at the present time one with observers, the measurement of $\hat{A}$ at the present time would select the eigenvalue $01$ of $\hat{A}$, namely “existence of sentient life” and the corresponding eigenvalue $01$ of the value of the fundamental physical constants observable $\hat{B}$. Since the universe’s evolution does not change the value of the fundamental constants, it is as if the eigenvalue of $\hat{B}$ was selected by a measurement of $\hat{B}$ at the beginning of time, when there could be no observers.
Of course, this is exactly the quantum search algorithm described in the previous section, the version with respect to the problem-solver of Table II. As shown by its time-symmetrization (Table IV), it can undergo with a quantum speedup (no matter the specific form of $\hat{\mathbf{U}}_{1,2}$) and does so even if the quantum search for the solution is random $\left[15\right]$.